1. a 1534 inch board is curt from a 36 14 inch board. The saw cut takes 38 inch. How much of the 36 14 inch board is left after cutting? Simplfiy your answer. type an interger simplified fraction, or mixed number. There are ___ inches left after cutting.

2. a dress is on sale for $176. what was the orginial price of the dress if the discount was 15 of the orginial price? the originial price of the dress was $____.

3.Al saves pennies. he agreed to give 37 of his pennies to bev is she would give 37 of what she got from al to carl and if carl in turn would give 37 of what he got from bev to dani, bev, carl and dani agreed and dani received 351 pennies. How many pennies did Al have intially? Al had _____pennies intially.
Answer #1

1.Given, Board size = 3614 inch

Size of board after cutting = 3614 - 1534 - 38 = 2042 inch

2. Price of dress after discount = $176

Let, original price of dress = x

Formula for calculating price after discount

Pd = { Po × ( 100 - d ) } ÷ 100

Here, Pd = price after discount = $176

Po = original price = x

d = discount = 15

Putting values in the formula we get

176 = { x × ( 100 - 15 ) } ÷ 100

x = $207.06

3. Let, Total number of pennies Al have = x

Number of pennies bev got from al = x × 0.37 = 0.37x

Number of pennies carl got from bev = 0.37 × 0.37x = 0.1369x

Number of pennies dani got from carl = 0.37 × 0.1369x = 0.050653x

Given, number of pennies Danny received = 351

So,

0.050653x = 351

x = number of pennies Al have = 351 ÷ 0.050653 = 6929.50 = 6930 ( approx )